# Gravitational Wave Spectral Shapes as a probe of Long Lived Right-handed Neutrinos, Leptogenesis and Dark Matter: Global versus Local BLB-L Cosmic Strings

### v1

@@ V.3 Co-genesis of Baryon Asymmetry and Dark Matter Asymmetry
In this section, we identify the parameter space that permits the simultaneous generation of the baryon asymmetry and the dark matter abundance with a period of matter domination. The entire framework ultimately traces its origin to the $U(1)_{B-L}$ symmetry. Its spontaneous breaking sets the scale for the right-handed neutrino masses, which govern both leptogenesis and dark matter production, and simultaneously generates the cosmic strings that source the stochastic gravitational wave background. The same mass scale determines when the right-handed neutrinos dominate the energy density, thereby modifying the expansion history and imprinting distinctive features on the gravitational wave spectrum. In this way, the right-handed neutrino masses, leptogenesis, dark matter, gravitational waves and the spectral shape modification are unified through their common origin in the $U(1)_{B-L}$ breaking scale. To consider this full framework and the co-genesis of baryon asymmetry and dark matter, we specify that the right-handed neutrino $N$ from the previous analysis is required to be the lightest right-handed neutrino $N_{1}$. To proceed with the analysis, it is convenient to introduce a new variable. To this end, we recall, Eq. 17, that the decay rate of a right-handed neutrino $N$ into leptons is
@@
which controls both the right-handed neutrino decay width and its connection to the light neutrino masses in the seesaw mechanism and the condition for matter domination. If the right-handed neutrino also couples to a dark matter candidate $\chi$ with a Yukawa coupling $y_{\rm DM}$, the additional decay channel $\Gamma_{N\to\chi}$ can be expressed in terms of an analogous parameter, which we denote the effective dark matter mass parameter,
@@
$\Gamma^{\rm tot}\;=\;\frac{M^{2}}{8\pi v^{2}}\,\tilde{m}_{\rm tot}\,,\quad\tilde{m}_{\rm tot}\;\equiv\;\tilde{m}+\tilde{m}_{\chi}\,.$ (93)
This parameterisation then allows every result in section II to be reproduced with the replacement $\tilde{m}\rightarrow\tilde{m}_{\rm tot}$. The condition for matter domination becomes,
$\tilde{m}_{tot}<1.1\times 10^{-8}\rm\ eV\ ,$ (94)
and the ending of matter domination occurs at,
$T_{\rm end}(M,\tilde{m}_{tot})\;\approx\;4.5\times 10^{-3}\,M\,\left(\frac{5.22\times 10^{-8}}{\tilde{m}_{tot}}-1\right)^{-0.68}$ (95)
@@
$\Delta=\left(1+3.726\times 10^{-6}\,\tilde{m}_{\rm tot}^{-0.67}\right)^{3/4}\ .$ (96)
@@
The dark matter mass, assuming it is less than the right-handed neutrino mass, is determined uniquely by the two effective mass parameters, which we show in figure 20.
@@
This demonstrates how a smaller effective dark matter mass parameter and a larger effective neutrino mass lead to a larger dark matter mass. From the contour analysis and imposing the bound $M_{\rm DM}>\mathcal{O}(10\rm KeV)$ from Lymann-$\alpha$ analysis, we find that viable dark matter requires $\tilde{m}_{\chi}\lesssim 3\times 10^{-10}$. This analysis can be straightforwardly extended to the case of a non-thermal initial abundance of right-handed neutrinos, where $Y_{N}^{i}$ in Eq. 99 is replaced by the corresponding non-thermal abundance specific to the model under consideration. An increased initial abundance enhances the resulting dark matter yield, thereby requiring a smaller dark matter mass to reproduce the observed relic density. The dark matter abundance can be expressed in two equivalent forms: one in terms of the effective masses and the initial abundance, and the other in terms of the dark matter mass, which can be written as follows
@@
$\simeq\frac{\tilde{m}_{\chi}}{\tilde{m}_{tot}}\ \left(1+3.726\times 10^{-6}\,\tilde{m}_{\rm tot}^{-0.67}\right)^{-3/4}\ Y_{N}^{i}\ .$
@@
We now turn to leptogenesis, which is determined by the lightest right-handed neutrino. Thus, we specify $N\rightarrow N_{1}$ and correspondingly set the mass as $M\rightarrow M_{1}$. In this case, the viable parameter space is determined by three quantities: the lightest right-handed neutrino mass $M_{1}$, the corresponding effective neutrino mass $\tilde{m}$, and the effective dark matter mass parameter $\tilde{m}_{\chi}$. Together, these parameters uniquely fix the region in which both the observed baryon asymmetry and the correct dark matter abundance can be obtained. We will show that successful dark matter production and baryogenesis can be readily achieved; however, since the testability of leptogenesis is limited to the boundary of the hierarchical regime, the detectability of both mechanisms remains severely constrained. In Fig. 21, we show the dependence of the baryon asymmetry on the effective masses. The scan ranges from the upper bound required for a period of matter domination down to the numerically determined minimum effective mass for successful baryogenesis, as obtained in Sec. IV.3. The baryon asymmetry grows with $\tilde{m}$ but decreases with $\tilde{m}_{\chi}$. This behaviour reflects the fact that the branching ratio into Standard Model particles increases with $\tilde{m}$ while it is suppressed by larger values of $\tilde{m}_{\chi}$.
@@
We performed a scan over the effective mass parameter space, bounded from below by the condition for successful leptogenesis and from above by the requirement of matter domination, while varying the right-handed neutrino mass. This allowed us to identify the regions compatible with successful leptogenesis. In Fig. 22, we illustrate the baryon asymmetry obtained for $M_{1}=10^{11}\mathrm{GeV}$ together with the corresponding bounds on the effective mass for different right-handed neutrino masses. The detectability in this scenario is highly limited, primarily due to the restricted testability of leptogenesis. As shown in Fig. 17, only a small region with $M_{1}<10^{10}\rm\ GeV$ lies within the range of potential experimental sensitivity.
Figure 22: Left panel: Scans of baryon asymmetry for $M_{1}=10^{11}$ GeV, the black dashed line denotes the observed baryon asymmetry, $Y_{B}^{\rm obs}$. Right Panel: $Y_{B}=Y_{B}^{\rm obs}$ for various right-handed neutrino masses. Successful leptogenesis is achieved if the parameters are to the right of the $Y_{B}^{\rm obs}$ lines. All points/lines on the plot have a period of matter domination in their evolution. A very narrow region of detectable parameter space appears at $M_{1}=10^{9}\rm GeV$ for both local and global cosmic strings; however, it lies at the extreme edge of the plotted range and is therefore not shown. For $M_{1}\geq 10^{10}\rm GeV$ there is no detectable signal. This illustrates that while leptogenesis and dark matter production can be simultaneously realised with relative ease, their experimental accessibility remains severely constrained by the limited detectability of leptogenesis itself.
@@ VI Discussion and Conclusion
In this work we have investigated the cosmological implications of long-lived heavy seesaw states within the framework of both global and gauged $U(1)_{B-L}$ symmetry breaking. Our analysis has identified a number of novel and testable features correlating right-handed neutrino dynamics, matter domination, gravitational waves, and the generation of the baryon asymmetry and dark matter.
@@
We first demonstrated that a period of early matter domination induced by heavy, long-lived particles can indeed occur. In the context of heavy seesaw states, right-handed neutrino domination arises in both the global and gauged $U(1)_{\rm B-L}$ realizations of the Type-I seesaw mechanism: the global case requires non-thermal production of right-handed neutrinos, whereas the gauged case can be achieved with purely thermal production. Crucially, we show that in the Type-II and Type-III seesaw frameworks, the heavy states cannot induce an early matter-dominated phase, as their interactions prevent an efficient freeze-out. These results hold independently of the light-neutrino mass ordering. By first deriving analytical estimates for the onset temperature and duration of the matter-dominated era, we established their dependence on the right-handed neutrino mass $M$ and the effective neutrino mass $\tilde{m}_{i}$. We then solved the full Boltzmann equations across all relevant parameter regimes and found that these analytical expectations are well captured by simple best-fit relations: the condition for matter domination reduces to a single requirement on the effective neutrino mass (Eq. 32), while the onset temperature and duration of the matter-dominated era are determined solely by $M$ (Eq. 31) and $\tilde{m}_{i}$ (Eq. 33), respectively.
@@
The key consequence of this early matter-dominated phase driven by right-handed neutrinos is its impact on the stochastic gravitational wave background generated by cosmic strings associated with $U(1)_{B-L}$ breaking. The modified expansion history alters the spectral shape of the GWB, providing a potential observational probe of RHN-induced matter domination. A particularly compelling and unique signature, for a transient, brief matter-dominated era due to these mesta-stable RHN, is a sharp, step-like feature in the GW spectrum, depicting observable kinks in the GW spectrum, one corresponding to the onset of matter domination and the other to its end (see Fig. 7). We analysed whether such characteristic features in the GW spectrum could be detected in next generation GW detectors such as LISA and ET. We showed that in the context with global cosmic strings, one will be able to probe right-handed neutrino masses across nine orders of magnitude, from $M\sim 10\,\mathrm{GeV}$ up to $M\sim 10^{9}\,\mathrm{GeV}$, with sensitivity down to effective masses of order $\tilde{m}\sim 10^{-9}\,\mathrm{eV}$. The same for local strings extend this reach considerably, covering right-handed neutrino masses from $M\sim 0.1\,\mathrm{GeV}$ scale to $10^{9}\,\mathrm{GeV}$, while probing neutrino mass parameters as small as $\tilde{m}\sim 10^{-10}\,\mathrm{eV}$. Symmetry breaking with a larger $v_{\rm B-L}$ leads to an increased amplitude of the gravitational-wave background and shifts its characteristic frequencies to lower values, and is therefore easier to detect the GW signal itself and the characteristic features on top of it (see Figs. 8 and 9). We presented some analytical results correlating the characteristic frequencies involving the break and knee and the seesaw parameters $\tilde{m}$, $M$ (see Eqns. (4351 and 31, 34).) via identifying the start and end temperatures of the period of RHN domination. The results are different for gauged or global BLB-L extensions. By estimating the SNR for various GW detectors ( 8 and 9) and combining different sets of GW detectors we found the discovery regions of such features which provide invaluable information and a concrete evidence for a new stage in the cosmological expansion history, enabling us to pin down the start and end of $N$ domination, thereby determining the suppression of the RHN mass scale ($M_{1}$) compared to the scale of spontaneous symmetry breaking ($v_{\rm B-L}$). The step-by-step pathway connecting measurable features of the gravitational-wave spectrum to the fundamental seesaw parameters is summarised in Fig. 11. The resulting novel features compatible with observed baryon asymmetry and DM relic, for instance, LISA will be able to detect $M_{1}\sim 10^{6}$ GeV. The detectable regions of the parameter space are shown in Figure 10, showing that a large region of the parameter space can be probed through gravitational-wave observations.
@@
Finally, we extended the analysis to scenarios where the right-handed neutrinos also decay into a dark sector. In this context, we examined both symmetric and asymmetric dark-matter production, deriving bounds on the viable dark-matter mass consistent with the observed relic density. The dark-matter abundance is likewise affected by entropy dilution from the late decays of the right-handed neutrinos, which modifies the relation between the decay parameters and the final relic density. We showed that the co-genesis of dark matter and the baryon asymmetry is straightforward, although its testability is primarily determined by the leptogenesis constraints. If leptogenesis is relaxed, however, near-future gravitational-wave and collider experiments could readily probe the region of parameter space where the heavy neutrinos predominantly decay into the dark sector. The relevant detectable parameter space for this scenario is shown in Figure 20. The mass for dark matter can range between the lymann alpha bound $\mathcal{O}(10\ \rm KeV)$ and the right-handed neutrino mass, $M$, which value is then fixed by the two effective neutrino mass parameters $\tilde{m},\ \tilde{m}_{\chi}$.
Once the characteristic features of the GW spectral shapes alluded to in this study are observed, one may look to target additional observations to distinguish between a metastable RHN-dominated pre-BBN era and other forms of early matter domination. Particularly in low-scale ARS leptogenesisAkhmedov_1998 ; Drewes_2018 ; Klaric_2021 RHN masses of GeV-TeV could be searched in typical Heavy Neutral Lepton Searches (HNL) search experiments (see Beacham:2019nyx ; Abdullahi:2022jlv ; Middleton:2022dio for current constraints and future experimental sensitivities) at the particle physics laboratories and astrophysical observables. In this manner, we can complement GW searches with laboratory searches in the same BSM parameter space. Nonetheless, such a study is beyond the scope of the present paper. We will explore this in future. Another complementary search for RH neutrino, if they exist, involves experiments such as neutrino-less double beta decay Dolinski:2019nrj ; Gomez-Cadenas:2011oep or lepton number violating processes Li:2021fvw ; Cai:2017mow and therefore provide us with a myriad of pathways to independently verify the existence of an early RHN-domination era, see Ref. Borboruah:2025hai for first steps towards such a complementary study but in the context with inflationary GW. This leads to a unique and exciting opportunity to form synergies between GW observations and laboratory searches.

### v2

@@ V.3 Co-genesis of Baryon Asymmetry and Dark Matter Asymmetry
In this section, we identify the parameter space that permits the simultaneous generation of the baryon asymmetry and the dark matter abundance with a period of early matter domination. The entire framework ultimately traces its origin to the $U(1)_{B-L}$ symmetry. Its spontaneous breaking sets the scale for the right-handed neutrino masses, which govern both leptogenesis and dark matter production, and simultaneously generates the cosmic strings that source the stochastic gravitational wave background. The same mass scale determines when the right-handed neutrinos dominate the energy density, thereby modifying the expansion history and imprinting distinctive features on the gravitational wave spectrum. In this way, the right-handed neutrino masses, leptogenesis, dark matter, gravitational waves and the spectral shape modification are unified through their common origin in the $U(1)_{B-L}$ breaking scale. To consider this full framework and the co-genesis of baryon asymmetry and dark matter, we specify that the right-handed neutrino $N$ from the previous analysis is required to be the lightest right-handed neutrino $N_{1}$. To proceed with the analysis, it is convenient to introduce a new variable. To this end, we recall, Eq. 17, that the decay rate of a right-handed neutrino $N$ into leptons is
@@
which controls both the right-handed neutrino decay width and its connection to the light neutrino masses in the seesaw mechanism and the condition for early matter domination. If the right-handed neutrino also couples to a dark matter candidate $\chi$ with a Yukawa coupling $y_{\rm DM}$, the additional decay channel $\Gamma_{N\to\chi}$ can be expressed in terms of an analogous parameter, which we denote the effective dark matter mass parameter,
@@
$\Gamma^{\rm tot}\;=\;\frac{M^{2}}{8\pi v_{H}^{2}}\,\tilde{m}_{\rm tot}\,,\quad\tilde{m}_{\rm tot}\;\equiv\;\tilde{m}+\tilde{m}_{\chi}\,.$ (99)
This parameterisation then allows every result in section II to be reproduced with the replacement $\tilde{m}\rightarrow\tilde{m}_{\rm tot}$. The condition for early matter domination becomes,
$\tilde{m}_{\rm tot}<1.1\times 10^{-8}\rm\ eV\ ,$ (100)
and the ending of early matter domination occurs at,
$T_{\rm end}(M,\tilde{m}_{\rm tot})\;\approx\;4.5\times 10^{-3}\,M\,\left(\frac{5.22\times 10^{-8}\ \rm eV}{\tilde{m}_{\rm tot}}-1\right)^{-0.68}$ (101)
@@
$\Delta=\left(1+3.726\times 10^{-6}\,\left(\frac{\tilde{m}_{\rm tot}}{\rm eV}\right)^{-0.67}\right)^{3/4}\ .$ (102)
@@
The dark matter mass, which has to be less than the right-handed neutrino mass, is determined uniquely by the two effective mass parameters, which we show in figure 20.
@@
This demonstrates how a smaller effective dark matter mass parameter and a larger effective neutrino mass lead to a larger dark matter mass. From the contour analysis and imposing the bound $M_{\rm DM}>\mathcal{O}(10\rm KeV)$ from Lyman-$\alpha$ analysis, we find that viable dark matter requires $\tilde{m}_{\chi}\lesssim 3\times 10^{-10}$. This analysis can be straightforwardly extended to the case of a non-thermal initial abundance of right-handed neutrinos, where $Y_{N}^{i}$ in Eq. 105 is replaced by the corresponding non-thermal abundance specific to the model under consideration. An increased initial abundance enhances the resulting dark matter yield, thereby requiring a smaller dark matter mass to reproduce the observed relic density. The dark matter abundance can be expressed in two equivalent forms: one in terms of the effective masses and the initial abundance, and the other in terms of the dark matter mass, which can be written as follows
@@
$\simeq\frac{\tilde{m}_{\chi}}{\tilde{m}_{tot}}\ \left(1+3.726\times 10^{-6}\,\left(\frac{\tilde{m}_{\rm tot}}{\rm eV}\right)^{-0.67}\right)^{-3/4}\ Y_{N}^{i}\ .$
@@
We now turn to leptogenesis, which is determined by the lightest right-handed neutrino. Thus, we specify $N\rightarrow N_{1}$ and correspondingly set the mass as $M\rightarrow M_{1}$. In this case, the viable parameter space is determined by three quantities: the lightest right-handed neutrino mass $M_{1}$, the corresponding effective neutrino mass $\tilde{m}$, and the effective dark matter mass parameter $\tilde{m}_{\chi}$. Together, these parameters uniquely fix the region in which both the observed baryon asymmetry and the correct dark matter abundance can be obtained. We will show that successful dark matter production and baryogenesis can be readily achieved; however, since the testability of leptogenesis is limited to the boundary of the hierarchical regime, the detectability of both mechanisms remains severely constrained. In Fig. 21, we show the dependence of the baryon asymmetry on the effective masses. The scan ranges from the upper bound required for a period of early matter domination down to the numerically determined minimum effective mass for successful baryogenesis, as obtained in Sec. IV.3. The baryon asymmetry grows with $\tilde{m}$ but decreases with $\tilde{m}_{\chi}$. This behaviour reflects the fact that the branching ratio into Standard Model particles increases with $\tilde{m}$ while it is suppressed by larger values of $\tilde{m}_{\chi}$.
@@
We performed a scan over the effective mass parameter space, bounded from below by the condition for successful leptogenesis and from above by the requirement of early matter domination, while varying the right-handed neutrino mass. This allowed us to identify the regions compatible with successful leptogenesis. In Fig. 22, we illustrate the baryon asymmetry obtained for $M_{1}=10^{11}\mathrm{GeV}$ together with the corresponding bounds on the effective mass for different right-handed neutrino masses. The detectability in this scenario is highly limited, primarily due to the restricted testability of leptogenesis. As shown in Fig. 17, only a small region with $M_{1}<10^{10}\rm\ GeV$ lies within the range of potential experimental sensitivity.
Figure 22: Left panel: Scans of baryon asymmetry for $M_{1}=10^{11}$ GeV, the black dashed line denotes the observed baryon asymmetry, $Y_{B}^{\rm obs}$. Right Panel: $Y_{B}=Y_{B}^{\rm obs}$ for various right-handed neutrino masses. Successful leptogenesis is achieved if the parameters are to the right of the $Y_{B}^{\rm obs}$ lines. All points/lines on the plot have a period of early matter domination in their evolution. A very narrow region of detectable parameter space appears at $M_{1}=10^{9}\rm GeV$ for both local and global cosmic strings; however, it lies at the extreme edge of the plotted range and is therefore not shown. For $M_{1}\geq 10^{10}\rm GeV$ there is no detectable signal. This illustrates that while leptogenesis and dark matter production can be simultaneously realised with relative ease, their experimental accessibility remains severely constrained by the limited detectability of leptogenesis itself.
@@ VI Discussion and Conclusion
In this work we have investigated the cosmological implications of long-lived heavy seesaw states within the framework of both global and gauged $U(1)_{B-L}$ symmetry breaking. Our analysis has identified a number of novel and testable features correlating right-handed neutrino dynamics, early matter domination, gravitational waves, and the generation of the baryon asymmetry and dark matter.
@@
We first demonstrated that a period of early matter domination induced by heavy, long-lived particles can indeed occur. In the context of heavy seesaw states, right-handed neutrino domination arises in both the global and gauged $U(1)_{\rm B-L}$ realizations of the Type-I seesaw mechanism: the global case requires non-thermal production of right-handed neutrinos, whereas the gauged case can be achieved with purely thermal production. Crucially, we show that in the Type-II and Type-III seesaw frameworks, the heavy states cannot induce an early matter-dominated phase, as their interactions prevent an efficient freeze-out. These results hold independently of the light-neutrino mass ordering. By first deriving analytical estimates for the onset temperature and duration of the matter-dominated era, we established their dependence on the right-handed neutrino mass $M$ and the effective neutrino mass $\tilde{m}_{i}$. We then solved the full Boltzmann equations across all relevant parameter regimes and found that these analytical expectations are well captured by simple best-fit relations: the condition for early matter domination reduces to a single requirement on the effective neutrino mass (Eq. 32), while the onset temperature and duration of the matter-dominated era are determined solely by $M$ (Eq. 31) and $\tilde{m}_{i}$ (Eq. 33), respectively.
@@
The key consequence of this early matter-dominated phase driven by right-handed neutrinos is its impact on the stochastic gravitational wave background generated by cosmic strings associated with $U(1)_{B-L}$ breaking. The modified expansion history alters the spectral shape of the GWB, providing a potential observational probe of RHN-induced early matter domination. A particularly compelling and unique signature, for a transient, brief matter-dominated era due to these mesta-stable RHN, is a sharp, step-like feature in the GW spectrum, depicting observable kinks in the GW spectrum, one corresponding to the onset of early matter domination and the other to its end (see Fig. 7). We analysed whether such characteristic features in the GW spectrum could be detected in next generation GW detectors such as LISA and ET. We showed that in the context with global cosmic strings, one will be able to probe right-handed neutrino masses across nine orders of magnitude, from $M\sim 10\,\mathrm{GeV}$ up to $M\sim 10^{9}\,\mathrm{GeV}$, with sensitivity down to effective masses of order $\tilde{m}\sim 10^{-9}\,\mathrm{eV}$. The same for local strings extend this reach considerably, covering right-handed neutrino masses from $M\sim 0.1\,\mathrm{GeV}$ scale to $10^{9}\,\mathrm{GeV}$, while probing neutrino mass parameters as small as $\tilde{m}\sim 10^{-10}\,\mathrm{eV}$. Symmetry breaking with a larger $v_{\rm B-L}$ leads to an increased amplitude of the gravitational-wave background and shifts its characteristic frequencies to lower values, and is therefore easier to detect the GW signal itself and the characteristic features on top of it (see Figs. 8 and 9). We presented some analytical results correlating the characteristic frequencies involving the break and knee and the seesaw parameters $\tilde{m}$, $M$ (see Eqns. (4957 and 31, 34).) via identifying the start and end temperatures of the period of RHN domination. The results are different for gauged or global BLB-L extensions. By estimating the SNR for various GW detectors ( 8 and 9) and combining different sets of GW detectors we found the discovery regions of such features which provide invaluable information and a concrete evidence for a new stage in the cosmological expansion history, enabling us to pin down the start and end of $N$ domination, thereby determining the suppression of the RHN mass scale ($M_{1}$) compared to the scale of spontaneous symmetry breaking ($v_{\rm B-L}$). The step-by-step pathway connecting measurable features of the gravitational-wave spectrum to the fundamental seesaw parameters is summarised in Fig. 11. The resulting novel features compatible with observed baryon asymmetry and DM relic, for instance, LISA will be able to detect $M_{1}\sim 10^{6}$ GeV. The detectable regions of the parameter space are shown in Figure 10, showing that a large region of the parameter space can be probed through gravitational-wave observations.
@@
Finally, we extended the analysis to scenarios where the right-handed neutrinos also decay into a dark sector. In this context, we examined both symmetric and asymmetric dark-matter production, deriving bounds on the viable dark-matter mass consistent with the observed relic density. The dark-matter abundance is likewise affected by entropy dilution from the late decays of the right-handed neutrinos, which modifies the relation between the decay parameters and the final relic density. We showed that the co-genesis of dark matter and the baryon asymmetry is straightforward, although its testability is primarily determined by the leptogenesis constraints. If leptogenesis is relaxed, however, near-future gravitational-wave and collider experiments could readily probe the region of parameter space where the heavy neutrinos predominantly decay into the dark sector. The relevant detectable parameter space for this scenario is shown in Figure 20. The mass for dark matter can range between the Lyman alpha bound $\mathcal{O}(10\ \rm KeV)$ and the right-handed neutrino mass, $M$, which value is then fixed by the two effective neutrino mass parameters $\tilde{m},\ \tilde{m}_{\chi}$.
Once the characteristic features of the GW spectral shapes alluded to in this study are observed, one may look to target additional observations to distinguish between a metastable RHN-dominated pre-BBN era and other forms of early matter domination. Particularly in low-scale ARS leptogenesisAkhmedov et al. (1998); Drewes et al. (2018); Klaric et al. (2021) RHN masses of GeV-TeV could be searched in typical Heavy Neutral Lepton Searches (HNL) search experiments (see Beacham and others (2020); Abdullahi and others (2023); Middleton (2022) for current constraints and future experimental sensitivities) at the particle physics laboratories and astrophysical observables. In this manner, we can complement GW searches with laboratory searches in the same BSM parameter space. Nonetheless, such a study is beyond the scope of the present paper. We will explore this in future. Another complementary search for RH neutrino, if they exist, involves experiments such as neutrino-less double beta decay Dolinski et al. (2019); Gomez-Cadenas et al. (2012) or lepton number violating processes Li et al. (2022); Cai et al. (2018) and therefore provide us with a myriad of pathways to independently verify the existence of an early RHN-domination era, see Ref. Borboruah et al. (2025) for first steps towards such a complementary study but in the context with inflationary GW. This leads to a unique and exciting opportunity to form synergies between GW observations and laboratory searches.